OpenStudy (anonymous):

Tan7pi/8 using half life formulas, find the exact value. Can someone help me solve this

zepdrix (zepdrix):

Hey c:

zepdrix (zepdrix):

Tangent Half Angle? Hmm, there are a bunch of different forms of that one, ya? How bout we use this one?\[\Large\rm \tan\left(\frac{x}{2}\right)=\frac{\sin x}{1+\cos x}\] So we setup our half angle first: \[\Large\rm \tan\left(\frac{7\pi}{8}\right)=\tan\left(\frac{7\pi/4}{2}\right)\]And then plug in the stuff,\[\Large\rm =\frac{\sin\left(\frac{7\pi}{4}\right)}{1+\cos\left(\frac{7\pi}{4}\right)}\]

OpenStudy (michele_laino):

do you know how to compute sin(7 pi/4) and cos(7 pi/4) ?

OpenStudy (michele_laino):

hint: \[\Large \begin{gathered} \sin \left( {\frac{{7\pi }}{4}} \right) = \sin \left( {2\pi - \frac{\pi }{4}} \right) = \hfill \\ = \sin \left( {2\pi } \right)\cos \left( {\frac{\pi }{4}} \right) - \cos \left( {2\pi } \right)\sin \left( {\frac{\pi }{4}} \right) = \hfill \\ = 0 \times \frac{1}{{\sqrt 2 }} - 1 \times \frac{1}{{\sqrt 2 }} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

\[\tan \alpha/2=\sqrt{(1-\cos)/(1+\cos)}\]

OpenStudy (anonymous):

This is the formula we used

OpenStudy (anonymous):

@zepdrix and @Michele_Laino

OpenStudy (michele_laino):

so, we can write: \[\Large \tan \left( {\frac{{7\pi }}{8}} \right) = \sqrt {\frac{{1 - \cos \left( {\frac{{7\pi }}{4}} \right)}}{{1 + \cos \left( {\frac{{7\pi }}{4}} \right)}}} \]

OpenStudy (anonymous):

Where does the 4 come from?

OpenStudy (michele_laino):

since if alpha/2= 7 pi/8, then alpha = 7 pi/4

OpenStudy (michele_laino):

please keep in mind that, the formula that you used, is: \[\Large \tan \left( {\frac{\alpha }{2}} \right) = \sqrt {\frac{{1 - \cos \alpha }}{{1 + \cos \alpha }}} \]

OpenStudy (anonymous):

So we only divide the bottom by 2?

OpenStudy (anonymous):

To get 7pi/4

OpenStudy (michele_laino):

we have to multiply (7 pi/8) by 2, in order to get (7 pi/4)

OpenStudy (anonymous):

Can you show me the work you did because I'm still confused

OpenStudy (michele_laino):

we have: \[\Large \begin{gathered} \cos \left( {\frac{{7\pi }}{4}} \right) = \cos \left( {2\pi - \frac{\pi }{4}} \right) = \hfill \\ = \cos \left( {2\pi } \right)\cos \left( {\frac{\pi }{4}} \right) + \sin \left( {2\pi } \right)\sin \left( {\frac{\pi }{4}} \right) = \hfill \\ = 1 \times \frac{1}{{\sqrt 2 }} + 0 \times \frac{1}{{\sqrt 2 }} = \frac{1}{{\sqrt 2 }} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

so, after substitution, we can write: \[\Large \tan \left( {\frac{{7\pi }}{8}} \right) = \sqrt {\frac{{1 - \frac{1}{{\sqrt 2 }}}}{{1 + \frac{1}{{\sqrt 2 }}}}} \]

OpenStudy (anonymous):

Got it, I'm going to use sqrt 2/2 instead of 1/sqrt2